# On the Natural Robustness of Vision-Language Models Against Visual Perception Attacks in Autonomous Driving

### v2

@@ 5 Discussion
Despite this progress, deploying VLMs on embedded AV hardware remains challenging due to limitations in compute power and energy efficiency. High-performance models such as LLaVA-7B and NVILA, though effective on server-grade GPUs, often require substantial memory and parallel processing capabilities that are impractical for in-vehicle deployment. One potential solution to reduce hardware demands is quantization—a widely used model compression technique for LLMs that improves computational efficiency by converting high-precision data types to lower-precision formats [23]. This process significantly reduces memory usage and model size, making it more feasible to run VLMs on edge devices. However, it may also introduce quantization errors, which could degrade precision [29]. We applied quantization to LLaVA-7B, expecting lower resource usage. Although initially expected to reduce latency, the quantization approach did not yield the desired outcome; instead, when tested on the NVIDIA A100 40 GB [1], it exhibited latencies ranging from 0.851 s to 2.158 s and 11.657 s at full precision, 8-bit, and 4-bit levels, respectively. Based on our preliminary analysis, we found that the main source of this issue could be due to an increase in demand stemming from the additional operations required to maintain accuracy in image processing tasks.
Another complementary strategy is adaptive image downsampling based on scene complexity. For example, in visually simple highway scenes, input images could be processed at lower resolutions to reduce computational load, while complex urban scenes could use higher resolutions to preserve important details. This approach creates a trade-off between efficiency and detection quality, as previously observed in object detection for AVs [7]. By adjusting the input resolution based on the type of scene, AV systems can improve resource management while maintaining reliable performance across different driving environments. In parallel, compression techniques such as model pruning and knowledge distillation (KD) have been successfully applied to reduce model size and inference cost in vision tasks. For instance, Rosa et al. [50] demonstrate that KD can yield a student model that is over 90% smaller than its teacher network while maintaining competitive accuracy in semantic segmentation. These techniques can complement downsampling strategies, further improving the feasibility of deploying VLMs on resource-constrained AV hardware.
6 Conclusion
In this work, we present Vehicle Vision Language Models (V2LMs) as a novel approach to enhance the robustness of AV perception systems against adversarial attacks. V2LMs outperform task-specific DNNs by maintaining high adversarial accuracy without requiring adversarial training, addressing a key limitation of traditional defense methods. The study evaluates two deployment strategies: Solo Mode, where separate VLMs are fine-tuned for common AV perception tasks such as TSR, ALC, and VD, as well as Tandem Mode, which uses a single unified V2LM for all three tasks, reducing storage requirements while maintaining comparable performance. Experimental results demonstrate that traditional DNN models suffer performance drops of 33%–46% under adversarial attacks, whereas V2LMs maintain significantly higher accuracy, with reductions of less than 8% on average. Additionally, this work explores the feasibility of integrating V2LMs into AV systems, proposing solutions such as model compression and adaptive image downsampling to mitigate latency challenges. This paper shows that V2LMs significantly enhance AV perception robustness, offering a promising path toward safer autonomous driving.

### v3

@@ V Discussion
This process significantly reduces memory usage and model size, making it more feasible to run VLMs on edge devices. However, it may also introduce quantization errors, which could degrade precision [24]. We applied quantization to LLaVA-7B, expecting lower resource usage. Although initially expected to reduce latency, the quantization approach did not yield the desired outcome; instead, when tested on the NVIDIA A100 40 GB, it exhibited latencies ranging from 0.851 s to 2.158 s and 11.657 s at full precision, 8-bit, and 4-bit levels, respectively. Based on our preliminary analysis, we found that the main source of this issue could be due to an increase in demand stemming from the additional operations required to maintain accuracy in image processing tasks.
VI Conclusion
We present V2LMs as a novel approach to enhance the robustness of AV perception systems against adversarial attacks. We evaluate Solo Mode and Tandem Mode, and demonstrate that V2LMs maintain high adversarial accuracy without adversarial training while reducing storage requirements and maintaining comparable performance. Experimental results show that task-specific DNNs suffer performance drops of 33%–74% under adversarial attacks, whereas V2LMs have reductions of less than 8% on average.
Acknowledgment
We gratefully acknowledge the support provided by the U.S. Department of Transportation (DOT) through the National Center for Transportation Cybersecurity and Resiliency (TraCR) under Grant No. 69A3552344812-2027534 and 69A3552348317. This work has also been partially supported by NSF under grant CNS-2443252 and The BMW Group. This research is with support from Google.org and the Google Cloud Research Credits program for the Gemma Academic Program.